# A clock with $8\times 10^{-19}$ systematic uncertainty

### v1

@@ Abstract
We report an optical lattice clock with a total systematic uncertainty of $8.1\times 10^{-19}$ in fractional frequency units, representing the lowest uncertainty of any clock to date. The clock relies on interrogating the ultra-narrow ${}^{1}S_{0}\rightarrow{}^{3}P_{0}$ transition in a dilute ensemble of fermionic strontium atoms trapped in a vertically-oriented, shallow, one-dimensional optical lattice. With $10^{5}$ atoms in Wannier-Stark eigenstates of this lattice, we measure record atomic coherence time and measurement precision reaching below $1\times 10^{-19}$ [1]. Such clock precision, together with imaging spectroscopy, enables precise control of collisional shifts as well as the lattice light shift [2, 3]. To address two remaining large systematic effects, we measure the second order Zeeman coefficient on the least magnetically sensitive clock transition, and we precisely determine the $5s4d$ ${}^{3}D_{1}$ lifetime to reduce the dynamic black body radiation shift uncertainty. All other systematic effects have uncertainties below $1\times 10^{-19}$.
@@
Introduction. Measuring time is one of the most fundamental tasks in physics, with each advancement in timekeeping enabling new discoveries and technologies [4]. Owing to the higher frequency of electronic transitions, the exceptional stability of optical metrology promises to revolutionize many disparate fields, from fundamental physics to navigation and geodesy. Over the past two decades, optical atomic clocks using neutral atoms or single ions have surpassed those based upon microwave transitions, setting records for both stability and accuracy [5, 6, 7]. Optical lattice clocks (OLCs) achieve ultra-high stability by simultaneously interrogating many atoms tightly confined within a standing wave of light. Every gain in stability and accuracy opens new realms of exploration, such as placing bounds on dark matter [8, 9], probing general relativity [10, 11], and will ultimately result in the redefinition of the SI second [12, 13, 14].
Building upon two decades of optical lattice clock development, the JILA 1D strontium OLC utilizes a shallow lattice formed within an in-vacuum build up cavity first described in Ref. [1]. The ${}^{87}$r ${}^{1}S_{0}\rightarrow{}^{3}P_{0}$ clock transition is addressed with a laser stabilized to a cryogenic, single-crystal silicon, optical resonator [15]. We previously reported record levels of atomic coherence and self-synchronous stability [1], cancellation of atomic interaction shifts [2], and precise control of the lattice light shift [3]. In this Letter, we report a complete systematic evaluation with a total uncertainty of $8.1\times 10^{-19}$ in fractional frequency units. Improved measurements of the second order Zeeman coefficient and the dynamic shift for black body radiation allow us to make significant strides in clock accuracy.
@@
High Accuracy Operation. To properly characterize and control systematic shifts, high measurement precision together with repeatable and reliable operation are key. Critical environmental control begins by stabilizing the air flow around the vacuum chamber to $20$ ${}^{\circ}$ with $100$ mK peak-to-peak stability. Each viewport flange on the vacuum chamber is temperature stabilized via separate liquid loops with better than $20$ mK stability. We directly measure the radiation temperature $T$ and its stability at the atom’s location by translating in a pair of calibrated in-vacuum temperature sensors [Fig. 1(a)].
An effusive oven generates a collimated beam of ${}^{87}$r that is slowed and cooled using the broad ${}^{1}S_{0}\rightarrow\,^{1}P_{1}$ transition at $461$ nm. The oven does not have direct line-of-sight to the main vacuum chamber, and we measure no temperature coupling to the oven at the atoms. This beam loads a magneto-optical trap (MOT) operating on this same transition. Further cooling on the ${}^{1}S_{0}\rightarrow\,^{3}P_{1}$ transition at $689$ nm reduces the temperature to a few $\mu$[16]. This cooling light is stabilized to the same silicon resonator as the clock laser with Hz-level drift per day, making cooling robust over many months. As shown in Fig. 1(a), an magic wavelength optical lattice at $813$ nm is formed within an in-vacuum buildup cavity oriented along gravity [1]. Atoms are loaded from the MOT into the lattice at a depth of $300$ lattice photon recoil energies ($E_{r}$). Doppler cooling on the ${}^{1}S_{0}\rightarrow{}^{3}P_{1}$ $F=11/2$ transition reduces the radial temperature to $\sim 700$ nK, and resolved motional sideband cooling along the tightly confined direction $\hat{Z}$ reduces the expected quantum state to $\langle n_{Z}\rangle<0.05$. During Doppler and sideband cooling, we polarize the the atomic sample in one of the $m_{F}=\pm 9/2$ stretched states.
@@
Radiation temperature. To ensure a fully thermal environment and measure the radiant temperature at the atoms, we follow a similar technique as in Refs. [18, 6]. Two calibrated thin film platinum resistance sensors are mounted to an in-vacuum translation arm, which is inserted into the middle of the vacuum chamber. This probe can be retracted to measure the temperature during operation. We observe a sub-mK temperature flicker floor at short time scales, $\sim 2$ mK peak oscillations on the hour time scale, and drift of less than a few mK per day, as shown in Fig. 1(d). These few-hour temperature fluctuations are from coupling to room temperature and building process chilled water and can likely be improved by further system isolation. At 12 hours the Allan deviation of the temperature is $1.4$ mK, which we treat as the operational stability. The total temperature uncertainty is $4.1$ mK [19].
@@
This precisely determined ${}^{3}D_{1}$ lifetime allows us to reevaluate $\nu_{dyn}$ based on the technique described in Ref. [22]. Using known atomic properties including measured transition strengths, magic wavelengths, and static polarizability, we determine $\nu_{dyn}=-153.06(33)$ mHz at 300 K [19]. The final BBR-related frequency shift combining both the static [23] and dynamic effects at the operational temperature of $20.132(4)$ ${}^{\circ}$ is $(-48417.2\pm 7.3)\times 10^{-19}$.
@@
[…]
@@
Density Shift. Although strong collisional shifts are suppressed by the fermionic nature of ${}^{87}$r, $p$-wave interactions lead to a systematic density shift [28, 29]. Previous Sr OLCs required separate evaluation of collisional shifts, leading to drifting systematics as sample preparation varied over time [6]. With imaging synchronously measuring different densities and frequency shifts throughout the sample, we perform real-time density shift corrections. As reported in Ref. [2], operating at a “magic lattice depth” near $15$ $E_{r}$, on-site $p$-wave and off-site $s$-wave interactions cancel each other, substantially reducing the density shift even with a large atomic sample. For a single characteristic run $\sim 300$ minutes, the correction is $(-1.1\pm 0.9)\times 10^{-19}$.
Other Systematic Shifts. Collisions between trapped strontium atoms and background gas result in a systematic frequency shift [30]. In this system, the background gas is dominated by hydrogen molecules [19]. As demonstrated in [31], the shift goes linearly with the vacuum lifetime. Using this coefficient and a measured vacuum lifetime of $63.6\pm 2.5$ s, we calculate a background gas shift of $(-4.7\pm 0.5)\times 10^{-19}$.
@@
Summary. Through precise atomic and environmental control, we have realized a strontium optical lattice clock with a total systematic accuracy of $8.1\times 10^{-19}$. This represents greater than a factor of $2$ improvement in systematic accuracy over the previously most accurate strontium optical lattice clock [6], and it sets the accuracy benchmark of all optical clocks reported to date. Black body radiation stands out as the most significant source of uncertainty, and future cryogenic operation should reduce uncertainty to the low $10^{-19}$ level [33].
Acknowledgements. We thank S. Porsev for accurate atomic structure calculations, D. Wellnitz and A. M. Rey for useful discussions on collective radiation effects, and C. Lisdat and U. Sterr for discussions about the dynamic BBR correction. We acknowledge contributions and discussions from the JILA Sr team, T. Bothwell, C. Kennedy, E. Oelker, J. Robinson, W. Tew, and A. Ellzey. We thank A. Chu and N. Darkwah Oppong for close reading of this manuscript. This research was supported in part through the use of University of Delaware HPC Caviness and DARWIN computing systems. Funding support is provided by NSF QLCI OMA-2016244, V. Bush Fellowship, NIST, and NSF Phys-1734006. K. K. is supported in part by the Quantum Information Research Support Center funded by NRF Korea, MSIT- No.2021M3H3A103657313.

### v2

@@ Abstract
We report an optical lattice clock with a total systematic uncertainty of $8.1\times 10^{-19}$ in fractional frequency units, representing the lowest uncertainty of any clock to date. The clock relies on interrogating the ultra-narrow ${}^{1}S_{0}\rightarrow{}^{3}P_{0}$ transition in a dilute ensemble of fermionic strontium atoms trapped in a vertically-oriented, shallow, one-dimensional optical lattice. Using imaging spectroscopy, we previously demonstrated record high atomic coherence time and measurement precision enabled by precise control of collisional shifts and the lattice light shift. In this work, we revise the black body radiation shift correction by evaluating the $5s4d$ ${}^{3}D_{1}$ lifetime, necessitating precise characterization and control of many body effects in the $5s4d$ ${}^{3}D_{1}$ decay. Lastly, we measure the second order Zeeman coefficient on the least magnetically sensitive clock transition. All other systematic effects have uncertainties below $1\times 10^{-19}$.
@@
Introduction. Measuring time is one of the most fundamental tasks in physics, with each advancement in timekeeping enabling new discoveries and technologies [1, 2]. Owing to the higher frequency of electronic transitions, the exceptional stability of optical metrology promises to revolutionize many disparate fields, from fundamental physics to navigation and geodesy. Over the past two decades, optical atomic clocks using neutral atoms or single ions have surpassed those based upon microwave transitions, setting records for both stability and accuracy [3, 4, 5]. Optical lattice clocks (OLCs) achieve ultra-high stability by simultaneously interrogating many atoms tightly confined within a standing wave of light [6, 7, 8]. Every gain in stability and accuracy opens new realms of exploration, such as placing bounds on dark matter [9, 10], probing general relativity [11, 12], and will ultimately result in the redefinition of the SI second [13, 14, 15].
Building upon two decades of optical lattice clock development, the JILA strontium 1D OLC utilizes a shallow lattice formed within an in-vacuum build up cavity first described in Ref. [16]. The 87Sr ${}^{1}S_{0}\rightarrow{}^{3}P_{0}$ clock transition is addressed with a laser stabilized to a cryogenic, single-crystal silicon, optical resonator [17]. We previously reported record levels of atomic coherence and self-synchronous stability [16], cancellation of atomic interaction shifts [18], and precise control of the lattice light shift [19]. In this Letter, we report a complete systematic evaluation with a total uncertainty of $8.1\times 10^{-19}$ in fractional frequency units. Improved measurements of the second order Zeeman coefficient and the dynamic shift for black body radiation allow us to make significant strides in clock accuracy.
@@
High Accuracy Operation. To properly characterize and control systematic shifts, high measurement precision together with repeatable and reliable operation are key. Critical environmental control begins by stabilizing the air flow around the vacuum chamber to $20$ C with $100$ mK peak-to-peak stability. Each viewport flange on the vacuum chamber is temperature stabilized via separate liquid loops with better than $20$ mK stability. We directly measure the radiation temperature $T$ and its stability at the atom’s location by translating in a pair of calibrated in-vacuum temperature sensors [Fig. 1(a)].
An effusive oven generates a collimated beam of 87Sr that is slowed and cooled using the broad ${}^{1}S_{0}\rightarrow\,^{1}P_{1}$ transition at $461$ nm. The oven does not have direct line-of-sight to the main vacuum chamber, and we measure no temperature coupling to the oven at the atoms. This beam loads a magneto-optical trap (MOT) operating on this same transition. Further cooling on the ${}^{1}S_{0}\rightarrow\,^{3}P_{1}$ transition at $689$ nm reduces the temperature to a few $\mu$[20]. This cooling light is stabilized to the same silicon resonator as the clock laser with Hz-level drift per day, making cooling robust over many months. As shown in Fig. 1(a), a magic wavelength optical lattice at $813$ nm is formed within an in-vacuum buildup cavity oriented along gravity [16]. Atoms are loaded from the MOT into the lattice at a depth of $300$ lattice photon recoil energies ($E_{r}$). Doppler cooling on the ${}^{1}S_{0}\rightarrow{}^{3}P_{1}$ $F=11/2$ transition reduces the radial temperature to $\sim 700$ nK, and resolved motional sideband cooling along the tightly confined direction $\hat{Z}$ reduces the expected quantum state to $\langle n_{Z}\rangle<0.05$. During Doppler and sideband cooling, we polarize the the atomic sample in one of the $m_{F}=\pm 9/2$ stretched states.
@@
Radiation temperature. To ensure a fully thermal environment and measure the radiant temperature at the atoms, we follow a similar technique as in Refs. [22, 4]. Two calibrated thin film platinum resistance sensors are mounted to an in-vacuum translation arm, which is inserted into the middle of the vacuum chamber. During clock spectroscopy the probe is retracted $30$ cm into an auxiliary vacuum chamber. We observe a sub-mK temperature flicker floor at short time scales, $\sim 2$ mK peak oscillations on the hour time scale, and drift of less than a few mK per day, as shown in Fig. 1(d). These few-hour temperature fluctuations are from coupling to room temperature and building process chilled water and can likely be improved by further system isolation. At 12 hours the Allan deviation of the temperature is $1.4$ mK, which we treat as the operational stability. The total temperature uncertainty is $4.1$ mK [23].
@@
This precisely determined ${}^{3}D_{1}$ lifetime allows us to reevaluate $\nu_{dyn}$ based on the technique described in Ref. [26]. Using known atomic properties including measured transition strengths, magic wavelengths, and static polarizability, we determine $\nu_{dyn}=-153.06(33)$ mHz at 300 K [23]. The final BBR-related frequency shift combining both the static [27] and dynamic effects at the operational temperature of $20.132(4)$ C is $(-48417.2\pm 7.3)\times 10^{-19}$.
@@
To determine $\xi_{\sigma\,m_{F}=5/2}$ precisely, we vary the applied bias field from $0.3$ to $1.5$ G and measure the resultant frequency shift in an interleaved manner, as shown in Fig. 3. $\Delta_{\text{vec}}$ is measured independently by modulating the lattice depth. We find a $\xi_{\sigma m_{F}=5/2}=(-0.12263\pm 0.00014)$ mHz/Hz2. At the operational field near $380$ mG, the second order Zeeman shift $\Delta\nu_{Z2}=(-85.51\pm 0.10)\times 10^{-18}$.
@@
Density Shift. Although strong collisional shifts are suppressed by the fermionic nature of 87Sr, $p$-wave interactions lead to a systematic density shift [32, 33]. Previous Sr OLCs required separate evaluation of collisional shifts, leading to drifting systematics as sample preparation varied over time [4]. With imaging synchronously measuring different densities and frequency shifts throughout the sample, we perform real-time density shift corrections. As reported in Ref. [18], operating at a “magic lattice depth” near $15$ $E_{r}$, on-site $p$-wave and off-site $s$-wave interactions cancel each other, substantially reducing the density shift even with a large atomic sample. For a single characteristic run $\sim 300$ minutes, the correction is $(-1.1\pm 0.9)\times 10^{-19}$.
Other Systematic Shifts. Collisions between trapped strontium atoms and background gas result in a systematic frequency shift [34]. In this system, the background gas is dominated by hydrogen molecules [23]. As demonstrated in [35], the shift is inversely proportional to the vacuum lifetime. Using this coefficient and a measured vacuum lifetime of $63.6\pm 2.5$ s, we calculate a background gas shift of $(-4.7\pm 0.5)\times 10^{-19}$.
@@
Summary. Through precise atomic and environmental control, we have realized a strontium optical lattice clock with a total systematic accuracy of $8.1\times 10^{-19}$ as reported in table 1. This represents greater than a factor of $2$ improvement in systematic accuracy over the previously most accurate strontium optical lattice clock [4], and it sets the accuracy benchmark of all optical clocks reported to date. Black body radiation stands out as the most significant source of uncertainty, and future cryogenic operation should reduce uncertainty to the low $10^{-19}$ level [37].
Acknowledgements. We thank S. Porsev for accurate atomic structure calculations, D. Wellnitz and A. M. Rey for useful discussions on collective radiation effects, and C. Lisdat and U. Sterr for discussions about the dynamic BBR correction. We acknowledge contributions and discussions from the JILA Sr team, T. Bothwell, C. Kennedy, E. Oelker, J. Robinson, W. Tew, and A. Ellzey. We thank A. Chu and N. Darkwah Oppong for close reading of this manuscript. This research was supported in part through the use of University of Delaware HPC Caviness and DARWIN computing systems. Funding support is provided by NSF QLCI OMA-2016244, V. Bush Fellowship, the DOE Quantum Systems Accelerator, NIST, and NSF PHY-2317149. K. K. is supported in part by the Quantum Information Research Support Center funded by NRF Korea, MSIT- No.2021M3H3A103657313.